OpenStudy (anonymous):

radical 1/2?? simplify

OpenStudy (lgbasallote):

\[\sqrt{\frac 12} \implies \frac{\sqrt 1}{\sqrt 2} \implies \frac{1}{\sqrt 2}\] now you can rationalize thatthing i last wrote right?

OpenStudy (anonymous):

no @lgbasallote ???

OpenStudy (anonymous):

@sasogeek ????

OpenStudy (anonymous):

But, technically you should not leave the square root in the denominator, so you would multiply by \[\sqrt{2} / \sqrt{2}\] and end up with: \[\frac{ \sqrt{2} }{ 2 }\]